Fermi National Laboratory


Accelerator Update

Monday October 27 through Wednesday October 29
All shifts began in shutdown.

Monday October 27
The day shift began with Booster in access. Beam was circulating in Booster by 1:00 PM. A transformer BW-2 tripped off halting Booster studies at 2:50 PM. Booster remained off once the transformer was reset in order to complete vacuum controls work.

Tuesday October 28
A Booster access was made on the Owl shift to investigate a high water temperature trip on GMPS (This is the power supply for the main Booster magnets), no problems were found. Beam was established to Booster during the day shift and studies began. Booster was turned off on the evening shift when a water leak was found in the Booster tunnel causing GMPS to trip off.

Wednesday October 29
Shutdown activities continue.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 10/29/2003   email Fermilab